### v1

# Versatile Demonstration Interface: Toward More Flexible Robot Demonstration Collection

Abstract
Previous methods for Learning from Demonstration leverage several approaches for a human to teach motions to a robot, including teleoperation, kinesthetic teaching, and natural demonstrations. However, little previous work has explored more general interfaces that allow for multiple demonstration types. Given the varied preferences of human demonstrators and task characteristics, a flexible tool that enables multiple demonstration types could be crucial for broader robot skill training. In this work, we propose Versatile Demonstration Interface (VDI), an attachment for collaborative robots that simplifies the collection of three common types of demonstrations. Designed for flexible deployment in industrial settings, our tool requires no additional instrumentation of the environment. Our prototype interface captures human demonstrations through a combination of vision, force sensing, and state tracking (e.g., through the robot proprioception or AprilTag tracking). Through a user study where we deployed our prototype VDI at a local manufacturing innovation center with manufacturing experts, we demonstrated the efficacy of our prototype in representative industrial tasks. Interactions from our study exposed a range of industrial use cases for VDI, clear relationships between demonstration preferences and task criteria, and insights for future tool design.
I Introduction
When teaching robots new skills, it can be difficult to effectively transfer expert knowledge to a robot’s behavior. Learning from Demonstration (LfD) [1] has emerged as a prevalent strategy for humans to teach such skills, where a person demonstrates a desired motion rather than relying on traditional low-level robot programming. Within LfD, several approaches have emerged for humans to provide demonstrations; including teleoperation, kinesthetic teaching, and natural interfaces [2]. Each of these methods has relative advantages and disadvantages that depend on a number of criteria, including task characteristics, individual demonstrator preferences, and the algorithm that encodes the robot skill. As an example, teleoperation is a convenient approach to collect demonstration data from a distance, but can suffer in terms of demonstration speed and quality depending on the teleoperation input device. On the other hand, natural demonstrations, such as video recordings of humans or instrumented tools, allow humans to show the robot exactly how they would perform the task. However, natural interfaces often suffer from correspondence issues between the human demonstrations and the robot task strategy (e.g., due to configuration limits). Given the varied characteristics and benefits of different demonstration modalities, we believe that future flexible automation systems [3] should make it easier to provide different types of demonstrations. In this work, we propose a demonstration interface (shown in Figure 1) that enables experts to provide demonstrations through three common demonstration modalities.
Figure 1: Versatile Demonstration Interface is a tool that connects to the end of a collaborative robot and makes it easier to collect demonstration data through teleoperation, kinesthetic teaching, and natural demonstrations (where the tool detaches from and is tracked by the robot).
Figure 2: Examples of the three demonstration modes of MMDI in a rolling task. (Left) The operator remotely controls the robot through a 6D input device. (Middle) The operator directly manipulates the robot arm and the environment forces are measured through a distal force sensor. (Right) The operator detaches the tool and performs demonstrations naturally while the robot optimizes its pose to track the tool with the interface’s camera.
With recent advancements in robot imitation learning, the need for new tools to collect demonstration data is pertinent. However, there has been little emphasis on creating general-purpose tools for collecting multiple forms of demonstrations. Ideally such general-purpose demonstration tools would be compatible with existing robot and task setups. Creating a single interface that combines demonstration modalities poses a number of fundamental challenges. This work addresses two key issues: how to design the interface for flexible use between demonstration modalities, and how to design the corresponding interaction modalities for data collection and to reduce the effort associated with switching between demonstration types.
This paper presents Versatile Demonstration Interface (VDI), a robot-mounted interface that enables users to provide three types of demonstrations during LfD. The primary contributions of this paper are the development of a prototype VDI interface and a user study with manufacturing experts that assesses its potential in manufacturing tasks. Our results suggest varied use cases for VDI and the corresponding interaction modalities. In the remaining sections, we review related work in the area of LfD, describe the development and features of our prototype, and present the findings from a user study where manufacturing experts performed a set of tabletop tasks using our prototype tool.
II Related Work
Our work develops a flexible LfD interface that combines multiple demonstration modalities. To contextualize our contributions, we provide a brief review of LfD methods and previous methods that assess and combine multiple demonstration modalities.
Learning from Demonstration Methods – From its introduction nearly thirty years ago [4], Learning from Demonstration (also often called programming by demonstration) has gained popularity as a paradigm to teach robots new skills, circumventing many challenges associated with traditional robot programming [1]. Past methods can generally can be categorized into three primary approaches – kinesthetic teaching, teleoperation, and passive observation (i.e., natural demonstrations) [5]. Kinesthetic teaching, where a robot arm is guided through training motions, is commercially available in many collaborative robot platforms and been used to teach robots a variety of complex skills [6, 7, 8, 9]. These methods are designed to be intuitive for end users and prioritize feasibility of skill transfer by inherently limiting the demonstrations to motions that can be achieved by the robot. The usability of such methods is often correlated with the control method, which frequently leverage gravity compensation, compliant, or impedance control [10, 11]. Teleoperation, where a robot is controlled from a distance, is a longstanding robot control paradigm [12] that can also be used to collect demonstrations [13]. More recently, vision-based learning systems have leveraged teleoperation to crowd source demonstration data that resembles deployment environments (i.e., with no human present) [14]. Natural demonstrations, where a person can perform a demonstration without specific considerations for robot execution, is a logical approach to collect high-quality and realistic human data, but poses many challenges such as correspondence (i.e., how to map human demonstrations to robot capabilities and limits) and state tracking [5]. To track the human, many recent methods leverage instrumented tools [15, 16, 17], augmented reality [18, 19], or learning from observations [20, 21].
With recent advances in imitation learning techniques [22, 23], interest in LfD paradigms that can collect data in an effective way has intensified. New physical interfaces have emerged to improve data collection through existing techniques. For example, the Universal Manipulation Interface (UMI) [24] is a low-cost natural demonstration interface for collecting prehensile demonstrations. Mobile ALOHA [25], HATO [26], and GELLO [27] are teleoperation setups to control bimanual robot arms with intuitive input devices across diverse household tasks.
Modality Comparisons and Fusing Modalities – Given the breadth of LfD approaches, previous work has systematically compared demonstration methods. Early comparisons highlighted the value of kinesthetic teaching over more traditional programming approaches (e.g., teach pendants) and teleoperated data collection [28, 29, 30, 31]. More recently, Maric et al. [32] quantified differences in performance and task load when performing contact-rich tasks using kinesthetic teaching and natural demonstrations and Praveena et al. [33] compared teleoperation, kinesthetic, hand, and instrumented tool demonstrations across various pick and place tasks. Generally, these previous studies suggest that participants perform better and prefer using more natural input methods (e.g., kinesthetic teaching over traditional programming, hand tracking input over kinesthetic teaching). However, in some cases, task characteristics may drive the choice of demonstration method. For example, a very precise task might benefit from teleoperation, where a user’s input can be scaled down (e.g., as is popular in robot-assisted surgery [34]). Alternatively, when operating a heavy tool, it may be beneficial for the operator to provide kinesthetic guidance over natural demonstrations to avoid the physical burden of the tool (e.g., through robot force offloading) [35]. When considering a flexible tool that can be used across many tasks, we thus see merit in a tool that combines demonstration modalities.
While limited work has investigated fusing together demonstration modalities, initial results suggest there may be benefits for both user interactions and policy learning. Open Multi-Purpose Interface (OMI) [36] combines kinesthetic teaching and teleoperation and exemplifies the benefits of a flexible interface in a multi-step home-care task using a mobile collaborative robot. In policy learning, MimicPlay [37] highlights the benefits of combining natural demonstrations with a small amount of teleoperated demonstrations to overcome robot correspondence issues. Similarly, PoCo [38] has shown how information from different demonstration modalities can be combined to learn more general skills.
III Versatile Demonstration Interface
To illustrate the opportunities associated with tools combining multiple demonstration modalities and to collect design and use case feedback, we developed an initial prototype of VDI. Here we discuss the important features of our physical and interaction design. Example interactions using our prototype interface are shown in the supplemental video.
III-A Design and Sensing
Our primary design goals were to (1) enable the three types of common demonstrations, (2) interface with existing end-of-arm tooling, and (3) require only sensing that is typically available on commercial robots or connected to the tool. In our prototype design, we also optimized for human factors when possible by following hand-tool design best practices. Our interface design schematic is shown in Figure 3.
[…]
To aid in LfD, the required sensing for data collection is included on-board the VDI interface. The interface consists of two main parts: the robot mount and the human interface. The robot mount includes a camera that can be used for visuomotor policy learning (i.e., collecting images of the tool during the demonstration) and for tracking the interface (e.g., the pose of the tool) when it is detached from the robot. The camera is placed in a fixed location with a known transform for localization to the robot base. For many tasks in industrial manufacturing, such as surface finishing, it is necessary for robots to regulate force in addition to kinematic goals. Thus, the tool includes a force sensor between the human interface and the end effector. This force sensor can be used to track applied forces when the human is grasping the interface. In our implementation, we use a uni-axial load cell sensor as a proof-of-concept force sensor (a 20 kg DYMH-103 Wheatstone bridge sensor with a HX711 amplifier), though such low-cost force sensors can suffer from nonlinearity and crosstalk effects [40]. We found that the resolution and repeatability was sufficient for our study, where the force information was primarily needed to identify mode switches. Future versions will employ a multi-axis sensor for higher fidelity data collection. A key design element is the ability to remove part of the interface from the robot to directly use it as a hand tool. To accommodate this feature, the robot mount includes a removable pin and force-sensitive resistor (FSR) that is used as a contact sensor. The human interface also employs a pattern of AprilTag [41] markers with known configurations that are tracked by the robot interface camera.
Our initial physical prototype had a few known design limitations that we plan to address in future versions. First, we included only a single camera that is used at times for data collection and at times for state tracking. For future visuomotor policy learning, it would beneficial to have an additional camera such that there is always a tool-centric view that can be used as part of the demonstration data. Our initial prototype also consists of a wired connection to a box containing the power electronics. We envision future iterations of the tool may include on-board electronics, batteries, wireless communication, and docked charging.
Figure 3: Schematic of the prototype interface. The interface consists of two main sections: the robot mount and the human interface. The schematic also illustrates the key features supporting the interaction modes.
III-B Demonstration Modalities
Teleoperation – In our implementation, the 6D teleoperation input to the robot is provided through a SpaceMouse, though our approach is amenable to any remote input device. Because the SpaceMouse does not provide force feedback to the operator, we operate our robot in an admittance control mode and use LEDs and sound to communicate when the level of applied force is approaching the set force limit. The gains (i.e., scaling) of the teleoperation input generally trade off precision with speed. In our implementation, we empirically tuned the gains for precise manipulations (as was required by the study tasks). The operator is able to switch into the teleoperation mode at any point when the tool is attached by pressing on the remote input device. If the operator exceeds the set force limit, the robot exits the teleoperation mode.
Kinesthetic Teaching – Our implementation leveraged the gravity compensation mode of the UR5e (i.e., free-drive mode) for kinesthetic interactions. We chose free-drive mode as a safe, passive solution for our user study, though depending on the collaborative robot platform, the kinesthetic teaching mode may also be achieved through impedance or admittance control. The tool-side force sensor is used to record interactions with the environment. The discrepancy between the robot force estimate and tool-side force estimate is used as a heuristic for whether to activate the kinesthetic mode. Intuitively, if the operator grabs onto and pulls on the interface, the robot will switch to the kinesthetic mode. In our implementation, the operator was required to pull in the tool direction because of the uni-axial force sensor. Depending on the force required for the task (and to kinesthetically guide the robot), the operator can choose to use either a power or precision grasp of the interface.
Natural Demonstrations – To engage in the natural demonstration mode, the operator can pull a pin on the robot mount and remove the human interface. Once the contact sensor detects that the tool is removed, the LEDs flash blue to indicate the system is ready to begin the natural mode. As soon as the camera locates one of the interface AprilTags, it will be begin to film the natural demonstration. To track the tool using the robot-mounted camera, we employ a nonlinear optimization of the robot pose that balances viewpoint heuristics, as is common in previous methods for shared control of robot cameras [42, 43, 44, 45, 46]. The optimization is a nonlinear weighted-sum of objectives subject to inequality constraints.
$\operatorname*{arg\,min}_{\textbf{p},\boldsymbol{\omega}}$ $\sum\limits_{i}w_{i}\phi_{i}(\textbf{p},\boldsymbol{\omega})$ (1)
s.t. $g_{i}(\textbf{p},\boldsymbol{\omega})\leq 0\;\forall i\in\{1,..,m\}$
where p is the commanded robot position, $\boldsymbol{\omega}$ is the robot orientation (in the form of a rotation vector), $w_{i}$ are the objective weights, $\phi_{i}$ are the objectives, and $g_{i}$ are the $m$ inequality constraints. In our implementation, the objectives select robot poses that (1) balance the distance between the camera and the tool, (2) a camera angle that centers the tool within the camera view and tracks the AprilTags from a specified angle, and (3) a small term that encourages a neutral (i.e., centered) camera pose and angle that we found empirically (through pilots) helps to avoid local minima in the optimization. The inequality constraints are used to limit the rotation and commanded position of the robot to within the robot’s dexterous workspace. As part of the optimization, we also limited the linear and angular velocity of the commanded robot pose for participant safety. The specific values and terms are in our open-source code [^1] https://github.com/mhagenow01/vdi. The state of the tool is estimated using the AprilTags. Each time a camera sees an AprilTag, the known configuration provides an estimate of the 6D tool pose. While we were able to design our prototype to include explicit state tracking, we are interested in the future to investigate other methods (or ensembles of approaches) for estimating the tool pose (e.g., FoundationPose [47], InfraredTags [48]). The various estimates of the tool pose are used as input to an Extended Kalman Filter (EKF) which provides the final pose estimation. In our implementation, the pose estimation was throttled to 5 Hz to allow for more observations per output pose estimate. For the purposes of our initial prototype implementation and optimizing the camera pose, we found this rate to be sufficient. However, for demonstration data collection, a higher estimation rate will be necessary. We imagine achieving a higher frequency through existing sensor fusion methods (e.g., adding an IMU to the interface). If the robot camera is unable to see any of the tool markers, the system beeps to alert the operator.
Figure 4: User study questionnaire results comparing the three demonstration modalities in terms of task load, usability, usefulness, and ease of use.
IV Evaluation
The primary purpose of our evaluation was to assess the feasibility of a tool that combines demonstration modalities and to assess how it might be applied in real-world manufacturing contexts. Thus, we aimed to collect usability data, design feedback, and to better understand potential use cases. The prototype was not designed for high-fidelity data collection, and thus, our experiment does not focus on data collection nor demonstrating the value of heterogeneous demonstration data in policy learning. We plan to assess these aspects in future studies with an updated interface.
IV-A Task Description
We designed two tasks to resemble processes in industrial manufacturing. Both task setups are shown in Figure 5. The first task was a rolling task, which proxies a manual thermomolding task and consisted of rolling a block of play-doh against a contoured mold. The play-doh was initially formed into a puck-like shape and secured to the mold using a pin (for one-handed completion). The purpose of the first task was to understand the perceived utility of the three demonstration modes in a complex, contact-rich task.
The second task was a press-fitting task where patterns of cylindrical fittings were pressed into a mold. The task was designed to combine coarse, fine, and force-controlled motions. Participants pressed sixteen fittings into a mold. The fittings were arranged in groups of eight with the two groups on opposite ends of the workbench. The purpose of the second study task was to further probe preferences for the demonstration modalities on a task with different characteristics (e.g., less path following, more coarse motion).
Figure 5: User evaluation tasks. Left: The press-fitting task where participants press sixteen fittings into designated slots. The two groups of fittings were on opposite ends of the workbench. Right: Rolling task where participants used the end effector to try and evenly distribute the material over the contoured surface. A pin helps to keep the material in place.
IV-B Evaluation with Manufacturing Experts
Our study involved nine experts (7M, 2F), aged 33–69 ($M=45.3,SD=11.1$), recruited from the Berkshire Innovation Center (BIC) via email. A summary of the participants is provided in Table I. Eight of the participants were from a manufacturing background with an average of 15 years of experience. The procedure was administered under a protocol #2402001226 approved by the Institutional Review Board (IRB) of the Massachusetts Institute of Technology. The experiment was designed to last an hour and a small remuneration was paid for participation.
ID Age Yrs Role Industry
P1 33 5 Maintenance Engineer Automotive
P2 57 30 Industrial Artist Sculptures
P3 34 3 Finance and Management Food Tech.
P4 39 20 Manager General Mfg
P5 42 24 Support Manager Paper
P6 52 16 Ortho. Massage Therapist Healthcare
P7 42 18 R&D Plastics
P8 40 1.5 Consultant General Mfg
P9 69 20 Plant Manager Chemicals
TABLE I: Summary of Participants
Procedure – Following informed consent, participants were briefed on the structure of the experiment. The participants then completed the rolling task using each of the conditions. The order of the conditions followed a Latin-square design to balance between controlling for order efforts and replicating order sequences. For each condition, the participant was briefed on how to use the method and given five minutes to interact with the method. After each condition, the participants filled out questionnaires.
After completing the three conditions of the rolling task, the participants completed the press-fitting task. Participants were instructed about the goals of the task and shown how to complete the task without the robot. Participants were prompted that they could switch demonstration modes at any point during the task, but that they needed to use each mode to complete at least one press fitting. Following the two tasks, participants engaged in a semi-structured interview where they were asked about their preferences, design feedback, and potential use cases for the prototype interface.
Metrics – After each condition of the rolling task, participants filled out the NASA Task Load Index (TLX) [49], the Systems Usability Scale (SUS) Questionnaire [50], and the Usefulness and Ease of Use 7-point subscales of the USE Questionnaire [51]. Given our small sample size, we report only descriptive statistics for the questionnaire responses. Responses from the semi-structured interview were synthesized into the key themes presented in the next section.
V Results and General Discussion
Here we discuss the results of our user study, limitations that we plan to address in future work, and general conclusions. All participants were generally able to complete the rolling task with each demonstration modality. In the press-fitting task, all participants were able to complete the task using the VDI tool. The questionnaire response results are shown in Figure 4 and participants ranked preferences are shown in Table II. Two participants responses are omitted from the questionnaire results due to missed survey items.
V-A User Study Interactions and Implications
The clear winner: natural demonstrations – The majority of participants ranked the natural mode as their top preference. Participants commented that it was the fastest way to accomplish the tasks [P7], was valuable for delicate motions [P4], and that they felt they could perform the task better compared to the other demonstration modes [P5, P8]. This was also reflected in the survey responses, where participants tended to rank the natural mode as lower task load and higher ease of use compared to the other two methods. Several participants attributed the benefits of this mode to its innate familiarity, including that it felt similar to using a pencil [P3] or a manual tool [P1]. For example, P2 commented “I am so used to working with my hands anyway, that it was a logical way to do things.” The most common complaint among participants, including those who didn’t prefer the natural mode, was the tracking accuracy. In the prototype, participants found that the range was very limited [P6] and that it could at times be difficult to guide the robot to certain views of the task. Given such limitations in range, P3 mentioned the natural mode may be most effective for 2D (e.g., tabletop) tasks where a similar camera angle can be used to track the hand tool longitudinally. Many participants commented that improvements to the tracking would increase their preference for the natural mode.
Method 1st 2nd 3rd Mean Rank
Teleoperation 0 4 5 2.56
Kinesthetic 3 4 2 1.89
Natural 6 1 2 1.56
TABLE II: Ranked preferences for demonstration methods
A time and place for each demonstration mode – While the natural mode was generally preferred, participants acknowledged the value of accessing the three demonstration modes through a single interface, both to cater to personal preferences and to task characteristics. As P7 commented, “With the engineers that I work with… we have a broad range of right-out-of-school to thirty years’ experience. People would pick different methods for sure.” Generally, the natural method was preferred for tasks involving complex and dexterous motions. For teleoperation, participants saw value for tasks where the demonstrator couldn’t be safely near the robot, such as environments with bio-hazards [P6] or heavy payloads [P7]. Some participants also commented on the value of using teleoperation to scale down human input for small, precise motions (e.g., soldering) [P4, P9] or to avoid the fatigue associated with more manual demonstration methods of repetitive tasks [P4]. For kinesthetic teaching, participants thought it would be valuable for tasks where more direct control was desired, but where the robot could assist with pushing/pulling [P5], or maintaining orientation [P1, P2]. P8 commented kinesthetic teaching was a good compromise between precision and ease of use “where I didn’t have to do too much, but I had some control.”
Participants also expressed some value in switching modes within tasks. Several participants mentioned using one mode for more coarse motions (e.g., kinesthetic teaching) and one mode for more precise motions (e.g., natural or teleoperation) [P2, P3, P9]. However, P1 commented that it may not be valuable to switch from teleoperation often if it is primarily used for tasks where the operator is removed from the robot.
Outlook and Improvements – Participants were generally optimistic about a range of repetitive tasks that could be trained using a versatile demonstration interface; including assembly (e.g., pick and place, pressing fittings, wire assembly) [P1, P4, P5, P7, P9], stoppering [P1], grinding [P2], quality measurements (e.g., coordinate measurements) [P4], process monitoring [P7], industrial food prep [P3], physical therapy [P6], and running research experiments [P7]. When probed about design improvements, most participants noted that they found the current prototype comfortable, including switching between multiple grasps for fine or coarse motor control [P8]. A few participants desired further flexibility to grasp the interface, such as through additional handles [P7] or modifications that accommodate two-handed grasps for better control [P9].
Participant suggestions and interactions identified several opportunities to improve the demonstration modes. For the natural mode, additional markers (particularly on the top of the device) may be beneficial for tasks where the camera is above the tool. Additionally, P3 suggested using a camera with a wider field of view. From observing participants, we also believe that enabling participants to manually grab and adjust the camera may be beneficial to escape undesired views by the robot optimization. For teleoperation, participants suggested that haptic feedback and other means of input might make teleoperation easier, particularly given the learning curve of the SpaceMouse [P9]. While participants generally thought it was simple to switch between the different modes, they suggested future design iterations could simplify the process of attaching and detaching the tool for the natural mode (e.g., requiring only one hand or prompting the robot to stop moving).
V-B Limitations and Future Work
In addition to the suggestions from our study participants, there were limitations of our setup and study that we plan to address in future work. We plan to test with several collaborative robots, different input devices for robot teleoperation, and through a wider suite of representative manufacturing tasks. Our study also only focused on perceptions of the VDI. Future iterations (with updated interface designs) will focus on characterizing the quality of data collection with the VDI modalities, investigating impacts of the multiple demonstration modalities on robot policy training, and exploring opportunities for robot to elicit targeted demonstration types based on task understanding.
V-C Conclusion
We presented Versatile Demonstration Interface (VDI), an interface that attaches to existing collaborative robot platforms and enables operators to teach robots through three types of demonstrations: teleoperation, kinesthetic, and natural. Through a user study with manufacturing experts, we identified opportunities and feedback for future interfaces toward flexible Learning from Demonstration.